OpenStudy (anonymous):

Solve r e^r=(1/2)^(t/80)

OpenStudy (klimenkov):

\[e^r=\left(\frac12\right)^\frac{t}{80}\]

OpenStudy (anonymous):

r = (-t/80) ln2

OpenStudy (anonymous):

Okay that sort of helps... what is the numerical value of r?

OpenStudy (anonymous):

if t be known.. r becomes specified... what is the t?

OpenStudy (anonymous):

Well it doesn't give me t so lets just say it's 1

OpenStudy (anonymous):

okey then... we know that ln2=0.69314718056 so... r = -1/80 * 0.69314718056 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got it right... thanks!